How do you find A.4

How do you find A.4
How do you find A.4
How do you find A.4
Answer #1

Part1: Preparation of Standard Solutions and Determination of Absorbances of the Standard Solutions: Preparation of Standard Solutions – known concentration of FeSCN2+

Fe3+    +    SCN-   à     [Fe(SCN)]2+

[FeSCN2+]/ [Fe3+][ SCN-] = Kc

In this set of beakers the concentration of Fe3+ ions is overwhelming greater than the concentration of SCN- ions. This shifts the equilibrium to the right so far that ALL the SCN- ions will react to become FeSCN2+ ions.

So in this set of beakers the concentration of the FeSCN2+ ions will be equal to the initial concentration of the SCN- ions. This initial concentration can be calculated using MV=MV to find the new concentration of ions in the new volume of 25.0mL.

With known concentration of FeSCN2+ we can calculate the absorbance of the solution using the measured percent absorbance. Graphing the Absorbance as a function of concentration will yield a line that will allow us to find the concentration of other solutions that we do not already know the concentration.

M1V1= M2V2  

0.001 (1ml) = x (25);   x = 0.00004 M

0.001 (2ml) = x (25);   x = 0.00008 M

0.001 (3ml) = x (25);   x = 0.00012 M

0.001 (4ml) = x (25);   x = 0.00016 M

0.001 (5ml) = x (25);   x = 0.00020 M

Initial concentration of SCN- ions and final concentration of FeSCN2+

Solution

Volume of 0.001M NaSCN (ml)

Initial concentration of SCN-

Final concentration of FeSCN2+ in 25 ml (Answer: A4)

1

0

0

0

2

1

0.00004 M

0.00004 M

3

2

0.00008 M

0.00008 M

4

3

0.00012 M

0.00012 M

5

4

0.00016 M

0.00016 M

6

5

0.00020 M

0.00020 M

From this known concentration of FeSCN2+ and absorbance value at lambda max we can plot a standard curve. Then by using equation for straight line we can calculate concentration of FeSCN2+ for any unknown sample. (Answer for A4 is last column in above table)